OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

x=-19

OpenStudy (texaschic101):

-4(x + 1) - 3 = -3(x - 4) -4x - 4 - 3 = -3x + 12 -4x - 7 = -3x + 12 -4x + 3x = 12 + 7 -x = 19 x = - 19